OpenStudy (anonymous):

Find the area of the region bounded by the graphs of y-1/x and 2x+2y=5

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (watchmath):

is that y=1/x ?

OpenStudy (anonymous):

yes. sorry

OpenStudy (watchmath):

The line 2x+2y=5 (i.e., y=(5-2x)/2) is above the graph of y=1/x and they intersect at (1/2,2) and (2,1/2) So the area is given by \[\int_{1/2}^2 (5-2x)/2 - \frac{1}{x}\,dx\] Try to continue from there. Watch this video to learn more about computing area http://www.youtube.com/watch?v=xWzycvfOpyw

OpenStudy (anonymous):

I think I have it, but what is the integral of x^-1

OpenStudy (watchmath):

ln x